# Transformers without Normalization

### v1

@@ 1 Introduction
Employing DyT is straightforward, as shown in Figure 1: we directly replace existing normalization layers with DyT in architectures such as vision and language Transformers. We empirically demonstrate that models with DyT can train stably and achieve high final performance across a wide range of settings. It often does not require tuning the training hyperparameters on the original architecture. Our work challenges the notion that normalization layers are indispensable for training modern neural networks and provides empirical insights into the properties of normalization layers. Moreover, preliminary measurements suggest that DyT improves training and inference speed, making it a candidate for efficiency-oriented network design.
@@ 4 Dynamic Tanh (DyT)
Integrating DyT layers into an existing architecture is straightforward: one DyT layer replaces one normalization layer (see Figure 1). This applies to normalization layers within attention blocks, FFN blocks, and the final normalization layer. Although DyT may look like or be considered an activation function, this study only uses it to replace normalization layers without altering any parts of the activation functions in the original architectures, such as GELU or ReLU. Other parts of the networks also remain intact. We also observe that there is little need to tune the hyperparameters used by the original architectures for DyT to perform well.
@@
DyT is not a new type of normalization layer, as it operates on each input element from a tensor independently during a forward pass without computing statistics or other types of aggregations. It does, however, preserve the effect of normalization layers in squashing the extreme values in a non-linear fashion while almost linearly transforming the very central parts of the input.
@@ 6 Analysis
We conduct several analyses on important properties of DyT. We begin by evaluating their computational efficiency, followed by two studies examining the roles of the tanh function and the learnable scale $\alpha$. Finally, we present comparisons with previous methods that aim to remove normalization layers.
6.1 Efficiency of DyT
We benchmark the LLaMA 7B model with RMSNorm or DyT by measuring the total time taken for 100 forward passes (inference) and 100 forward-backward passes (training) using a single sequence of 4096 tokens. Table 7 reports the time required for all RMSNorm or DyT layers and the entire model when running on an Nvidia H100 GPU with BF16 precision. DyT layers significantly reduce computation time compared to RMSNorm layers, with a similar trend observed under FP32 precision. DyT may be a promising choice for efficiency-oriented network design.
inference training
LLaMA 7B layer model layer model
RMSNorm 2.1s 14.1s 8.3s 42.6s
DyT 1.0s 13.0s 4.8s 39.1s
[…]
Table 7: Inference and training latency (BF16 precision) for LLaMA 7B with RMSNorm or DyT. DyT achieves a substantial reduction in both inference and training time.
@@ 8 Related Work
Many studies have explored how to train deep models without normalization layers. Several works (Zhang et al., 2019; De and Smith, 2020; Bachlechner et al., 2021) explore alternative weight initialization schemes to stabilize training. The pioneering work by Brock et al. (2021a, b) show that high-performing ResNets can be trained without normalization (Smith et al., 2023) through combination of initialization techniques (De and Smith, 2020), weight normalization (Salimans and Kingma, 2016; Huang et al., 2017; Qiao et al., 2019), and adaptive gradient clipping (Brock et al., 2021b). Additionally, their training strategy incorporates extensive data augmentation (Cubuk et al., 2020) and regularization (Srivastava et al., 2014; Huang et al., 2016). The studies above are based on various ConvNet models.
In Transformer architectures, He and Hofmann (2023) explore modifications to Transformer blocks that reduce reliance on normalization layers and skip connections. Alternatively, Heimersheim (2024) propose a method to gradually remove LN from pretrained networks by fine-tuning the model after removing each normalization layer. Unlike previous approaches, DyT requires minimal modifications to both the architecture and the training recipe. Despite its simplicity, DyT achieves stable training and comparable performance.
@@ 9 Limitations
We conduct experiments on networks using either LN or RMSNorm because of their popularity in Transformers and other modern architectures. Preliminary experiments (see Appendix C) indicate that DyT struggles to replace BN directly in classic networks like ResNets. It remains to be studied in more depth whether and how DyT can adapt to models with other types of normalization layers.
@@ Appendix B Hyperparameters
Appendix C Replacing Batch Normalization with DyT

### v2

@@ 1 Introduction
Employing DyT is straightforward, as shown in Figure 1: we directly replace existing normalization layers with DyT in architectures such as vision and language Transformers. We empirically demonstrate that models with DyT can train stably and achieve high final performance across a wide range of settings. It often does not require tuning the training hyperparameters on the original architecture. Our work challenges the notion that normalization layers are indispensable for training modern neural networks and provides empirical insights into the properties of normalization layers.
@@ 4 Dynamic Tanh (DyT)
Integrating DyT layers into an existing architecture is straightforward: one DyT layer replaces one normalization layer (see Figure 1). This applies to normalization layers within attention blocks, FFN blocks, and the final normalization layer. Although DyT may look like or be considered an activation function, this study only uses it to replace normalization layers without altering any parts of the activation functions in the original architectures, such as GELU or ReLU. Readers interested in the use of harmonic or hyperbolic functions as activation functions can refer to Hashemi et al. (2024). We also observe that there is little need to tune the hyperparameters used by the original architectures for DyT to perform well.
@@
DyT is not a new type of normalization layer, as it operates on each input element from a tensor independently during a forward pass without computing statistics or other types of aggregation. It does, however, preserve the effect of normalization layers in squashing the extreme values in a non-linear fashion while almost linearly transforming the very central parts of the input.
@@ 6 Analysis
In this section, we begin with ablations on the effects of the tanh function and the learnable scalar $\alpha$. We then analyze the values of $\alpha$ throughout and after training. Lastly, we present comparisons with previous methods that aim to remove normalization layers.
@@ 8 Related Work
Many studies have explored how to train deep models without normalization layers. Klambauer et al. (2017) introduce an alternative activation function that enables self-normalizing behavior, eliminating the need for explicit normalization. Other works (Zhang et al., 2019; De and Smith, 2020; Bachlechner et al., 2021) propose specialized initialization schemes to stabilize training in the absence of normalization. The pioneering work by Brock et al. (2021a, b) show that high-performing ResNets can be trained without normalization (Smith et al., 2023) through combination of initialization techniques (De and Smith, 2020), weight normalization (Salimans and Kingma, 2016; Huang et al., 2017; Qiao et al., 2019), and adaptive gradient clipping (Brock et al., 2021b). Additionally, their training strategy incorporates extensive data augmentation (Cubuk et al., 2020) and regularization (Srivastava et al., 2014; Huang et al., 2016). The studies above are based on various ConvNet models.
In Transformer architectures, He and Hofmann (2023) explore modifications to Transformer blocks that reduce reliance on normalization layers and skip connections. Jha and Reagen (2024) introduce AERO, a Softmax-only LLM that improves inference efficiency and privacy with minimal performance loss. Alternatively, Heimersheim (2024) propose a method to gradually remove LN from pretrained networks by fine-tuning the model after removing each normalization layer. Unlike previous approaches, DyT requires minimal modifications to both the architecture and the training recipe. Despite its simplicity, DyT achieves stable training and comparable performance.
@@ 9 Limitations
Our experiments focus on networks using LN or RMSNorm because of their popularity in Transformers and other modern architectures. Preliminary experiments (see Appendix D) indicate that DyT struggles to replace BN directly in classic networks like ResNets. It remains to be studied in more depth whether and how DyT can adapt to models with other types of normalization layers.
Furthermore, although DyT is conceptually and computationally simpler, we find that DyT offers no speedup over models with normalization layers when properly compiled/optimized (see Appendix C). Its computational benefits across different hardware platforms or deployment environments remain uncertain.
@@ Appendix B Hyperparameters
Appendix C Efficiency of DyT
We benchmark the LLaMA 7B model with RMSNorm or DyT by measuring the total time required for 100 forward passes (inference) and 100 forward-backward passes (training) on a single sequence of 4096 tokens. We first follow the officially recommended LLaMA setup and load the model from Hugging Face without applying any performance optimizations. Table 14 reports the time taken for RMSNorm and DyT layers, as well as for the entire model, when running on a Nvidia H100 GPU with BF16 precision. DyT layers reduce computation time compared to RMSNorm layers.
inference training
LLaMA 7B layer model layer model
RMSNorm 2.1s 14.1s 8.3s 42.6s
DyT 1.0s 13.0s 4.8s 39.1s
[…]
Table 14: Inference and training latency (BF16 precision) for LLaMA 7B with RMSNorm or DyT. DyT achieves a substantial reduction in both inference and training time. Results are measured without any extra performance optimizations.
We also benchmark both models using torch.compile. Interestingly, compiling the entire LLaMA model increases latency for the Hugging Face implementation, and compiling only the DyT or RMSNorm layers yields more efficient execution. Table 15 shows that, after compilation, the latency of the RMSNorm and DyT layers becomes nearly identical.
inference training
LLaMA 7B layer model layer model
RMSNorm 0.3s 12.3s 3.9s 38.9s
DyT 0.3s 12.3s 3.9s 38.9s
Table 15: Inference and training latency (BF16 precision) for a LLaMA 7B with compiled RMSNorm or DyT. After compilation, the latency of RMSNorm and DyT layers are nearly identical.
An important distinction of DyT is that it is an element-wise operation and does not require a reduction operation within itself, compared to normalization layers. This could make it faster on hardware where reduction is a bottleneck. Additionally, even on conventional GPUs, DyT could offer opportunities for further optimization, e.g., fusing it with the preceding matrix multiplication layer from the last residual block.
Appendix D Replacing Batch Normalization with DyT